OpenStudy (anonymous):

Simplify. 3^2 · 3^4 · 3^6

sammixboo (sammixboo):

First do the exponents, what is 3^2 or 3 x 3?

OpenStudy (anonymous):

\[x^a * x^b*x^c = x^{a+b+c}\]

OpenStudy (anonymous):

9

sammixboo (sammixboo):

Now we have 9 x 3^4 x 3^6 What is 3^4 or 3 x 3 x 3?

sammixboo (sammixboo):

3 x 3 x 3 x 3* I meant >_<

OpenStudy (anonymous):

\[3^2*3^4 *3^6 = 3^{(2+4+6)}\] \[3^{(2+4+6)} = 3^{12}\]

OpenStudy (anonymous):

thxs yall